OpenStudy (carolina4567):

Can somebody Check These Answers For ME :) Your friend hands you a graph of the performance statistics of the latest powerboat to be produced. He says, “I know this graph is f(x) = –2(x – 4)3 – 3 but I can’t remember how it is related to the graph of x3.” Explain to your friend how the graph f(x) is a translation of the graph x3. My answer - Well when you translate ( x^3 ) downwards you get f(x) = -2(x-4)^3-3. Meaning that f(x) = -2(x-4)^3-3 is a translation of ( x^3 ).

OpenStudy (carolina4567):

@iGreen @amistre64 @mathmale

OpenStudy (amistre64):

Your friend hands you a graph of the performance statistics of the latest powerboat to be produced. He says, “I know this graph is f(x) = –2(x – 4)^3 – 3 but I can’t remember how it is related to the graph of x^3.” Explain to your friend how the graph f(x) is a translation of the graph x3. You answered - Well when you translate ( x^3 ) downwards you get f(x) = -2(x-4)^3-3. Meaning that f(x) = -2(x-4)^3-3 is a translation of ( x^3 ). this is not accurate. there are 2 translations involved, and a reflection, and a whatever you call a stretching

OpenStudy (amistre64):

refresh to rid the odd marks ...

OpenStudy (amistre64):

if we start with a function f(x) we can see that it is shifted (translated side to side) by h units f(x-h) now we also see that it has been stretched by a, and reflected about the x axis: -a f(x-h) it has then been shifted (translated up/down) by k units: -a f(x-h) + k

OpenStudy (amistre64):

and since f(x) = x^3 .. then f(x-h) = (x-h)^3